OpenStudy (anonymous):

what is 6/5 as a percent

HanAkoSolo (jamierox4ev3r):

with problems like this, it's helpful to set it as a fraction over 100

OpenStudy (anonymous):

120%

HanAkoSolo (jamierox4ev3r):

@0278700km would you like an explanation for that answer?

OpenStudy (anonymous):

yes

HanAkoSolo (jamierox4ev3r):

alright :) So 5*20=100

HanAkoSolo (jamierox4ev3r):

if we have \(\Large\frac{6}{5}\), then we must multiply both denominator and numerator by 20

HanAkoSolo (jamierox4ev3r):

if we do that, then we have \(\Large\frac{120}{100}\) Since this fraction is over 100, and the number over it is 120, we know immediately that the percentage is 120%. While @Seedur was correct, he shouldn't have gave you the answer correctly, as it is against our rules here. I hope you have a better understanding of this problem :)

OpenStudy (anonymous):

\[\frac{6}{5}=1.2 \]Move the decimal point two places to the right and then append the % sign.

HanAkoSolo (jamierox4ev3r):

^ this is an especially handy trick to user, especially when your denominator doesn't divide perfectly into 100.

hero (hero):

Answer givers are starting to really anger me.

HanAkoSolo (jamierox4ev3r):

\(\Large\frac{6}{5}\)=1.2 ---> 120.0 --> 120%

HanAkoSolo (jamierox4ev3r):

@Hero unfortunately, they're becoming increasingly common. It's frustrating for me too

HanAkoSolo (jamierox4ev3r):

to have to give users private talks, and then at times, the users refuse to listen. As a moderator, you still have to deal with this much more than I do, so I sincerely respect the work that you and the rest of the mod. team do here on OS. :)